OpenStudy (anonymous):

4/6 < 2/3 ? 4/6 > 2/3 ? or 4/6 = 2/3

OpenStudy (hoblos):

4/6 = 2/3

OpenStudy (ash2326):

4/6 =2/3 so third option

OpenStudy (anonymous):

Thank You So Muchh (:

OpenStudy (hoblos):

my pleasure XD